Circular No. 9084

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2009 SK_{280} (SPACEWATCH-HILL) R. E. Hill, Lunar and Planetary Laboratory, reports his discovery of

a comet with the Mount Lemmon Survey’s 1.5-m f/2 reflector on Oct. 15.4 UT. Four co-added 30-s exposures showed a small, condensed nuclear condensation with a 20′′–30′′ fan-shaped tail in p.a. 240^{o}; the comet’s appearance was similar on four further co-added exposures an hour later. Pre-discovery Spacewatch observations (all showing asteroidal appearance) obtained by T. H. Bressi, J. V. Scotti, and R. S. McMillan on Sept. 17, 25, and 29 were then identified at the Minor Planet Center, those on the last two nights having previously been linked under the designation 2009 SK_{280} and published on MPS 297945. The two discovery observations are given below (the Spacewatch data from Sept. 25 being made with the 0.9-m f/3 reflector at Kitt Peak).

2009 UT α_{2000} δ_{2000} Mag. Observer
Sept.25.28819 1^{h}44^{m}47.^{s}49 +4^{o}15′40′′.7 20.1 Spacewatch
Oct. 15.36043 1 32 40.70 +3 54 44.8 19.8 Hill

On Oct. 16.3, Scotti (using the Spacewatch II 1.8-m f/2.7 reflector) detected a 0′.39 tail in p.a. 239^{o}, while Hill (at Mount Lemmon) noted a 20′′–30′′ tail in p.a. 230^{o} and W. H. Ryan (Magdalena Ridge, 2.4-m f/8.9 reflector) found a tail in p.a. 225^{o}.

The previously unpublished astrometric observations appear on MPEC 2009-U02, together with these elliptical orbital elements and an ephemeris:

T = 2009 Apr. 22.2398 TT  ω = 324.^{o}1478
e = 0.127467  Ω = 36.4871 }2000.0
q = 4.178926 AU  i = 16.7109
a = 4.789418 AU n^{o} = 0.0940329 P = 10.5 years

COMETS C/2009 J1 AND C/2009 J2 (SOHO) Further to IAUC 9083, additional Kreutz sungrazers have been found

on SOHO website images. C/2009 J1 was slightly diffuse with a short tail (peak mag ∼ 6.5–7). C/2009 J2 was very faint (mag ∼ 8–8.5) and diffuse.

Comet 2009 UT α_{2000} δ_{2000} Inst. F MPEC
C/2009 J1 May 3.008 2^{h}45.^{m}9 +14^{o}20′ C2 MU 2009-N46
C/2009 J2  3.458 2 47.6 +14 27 C2 MK 2009-N46
2009 October 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green